Find mole fraction of solute in 0.5 molal NaOH solution ? Sir, here how to find moles of solute.

Grade:11

1 Answers

Arun
3 years ago
Molatity = x .1000/(1 -x) M
where x is mole fraction, M is molar mass of solvent ,
now,
2 = 1000x/(1-x)18
 
36/1000 =x/(1-x)
 
0.036- 0.036x = x
 
0.036 =1.036x
 
x =0.036/1.036
x =0.03474